Measure of a complementary angle is 30 degrees more than twice the given angle. Find measure of complementry angle
Answer by LinnW(1048)

Set x = the measure of the angle.
The complement of x is 90-x

complement = 30 plus twice the angle
(90-x) = 30 + 2*x
add x to each side
90 = 30 + 3x
add -30 to each side
60 = 3x
divide each side by 3
20 = x
Since the complement of x = (90-x)
The complement is 90-x = 90-20 = 70